# SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant

### v1

@@ 4.3 Visual Information Discovery
Table 3: Ablation study of training strategy on visual instruction tasks. All models are in 7B scale with three components of ViT-LoRA (V-LoRA), self-questioning (SQ), and prototype extractor (Proto). We provide experiments on two instruction datasets Liu et al. (2023b); Chen et al. (2023) with different pre-training (PT) and instruction tuning (IT) data scales.
[…]
@@
4.4 Ablation Study
In Table 3, we conduct experiments with different architecture designs and training strategies on five question-answering benchmarks. For a fair comparison, we train the baseline models on our local machine with the same training recipe of LLaVA-LoRA Liu et al. (2023b). Specifically, we present our full model and three ablated models by removing one component each time, where all models are in 7B scale. We adopt the dataset Liu et al. (2023b) with $558$k for pre-training (PT) and $665$k for fine-tuning (FT). As compared, self-questioning (SQ) brings a consistent performance boost on all the five benchmarks, indicating the effectiveness of visual self-questioning on improving visual language understanding. Besides, we introduce the prototype extractor (Proto) to enhance visual representation, achieving $0.9\%$ improvement in average accuracy among five benchmark. With all three components incorporated, we observe a $2.4\%$ improvement in average accuracy.
As shown by the bottom block of Table 3, we conduct experiments with the same ablation settings but with a larger scale of the visual instruction data Chen et al. (2023) (i.e., for both PT and IT). Overall, SQ-LLaVA achieves $2.4\%$ improvement over the baseline model after training on the smaller dataset and achieves $3.0\%$ improvement after training on the larger dataset. We also provide ablation study of SQ-LLaVA in 13B scale (see Appendix A.1 for detail).
5 Conclusion
@@
This work introduced SQ-LLaVA, a method to improve general-purpose vision language understanding and image-oriented question answering through the proposed visual self-questioning technique. Our experiments reveal that SQ-LLaVA achieved superior performance with fewer training parameters. We performed a comprehensive study on visual instruction tuning tasks and found that SQ-LLaVA can generalize well to a wide range of unseen tasks and outperform state-of-the-art methods. Notably, qualitative assessments demonstrated that SQ-LLaVA allows enhanced visual representation and domain alignment capability, effectively alleviating the issue of object hallucination and improving the semantic interpretation of images. Our findings underscored the potential of visual self-questioning as a potent training strategy within the field of visual instruction tuning, paving the way for more efficient and effective model training approaches. SQ-LLaVA is a foundational step toward enhancing general vision language understanding by framing question-asking as an intrinsic goal of the question-answering process. We hope the proposed SQ-LLaVA can encourage future endeavors in exploiting questions in visual instruction tuning by following the proposed self-questioning technique.
Appendix A Zero-shot Multilingual Capability
A.1 Ablation Study
In Table 4, we conduct experiments with different architecture designs and training strategies on five benchmarks for SQ-LLaVA-13B. For a fair comparison, we train the baseline models on our local machine with the same training recipe of LLaVA-LoRA Liu et al. (2023b). Specifically, we present the baseline model, our full model, and three ablated models by removing one component each time. We adopt the dataset Liu et al. (2023b) with $558$k for pre-training (PT) and $665$k for fine-tuning (FT). As compared, self-questioning (SQ) achieves $0.6\%$ improvement in average accuracy, indicating the effectiveness of visual self-questioning in improving visual language understanding. Besides, we introduce the prototype extractor (Proto) to enhance visual representation, achieving $0.6\%$ improvement in average accuracy. With all three components incorporated, we observe a $2.5\%$ improvement in average accuracy among five benchmarks.
As shown by the bottom block of Table 4, we conduct experiments with the same ablation settings but with a larger scale of the visual instruction data Chen et al. (2023) (i.e., for both PT and IT). Overall, SQ-LLaVA achieves $2.5\%$ improvement over the baseline model after training on the smaller dataset and achieves $3.2\%$ improvement after training on the larger dataset.
In Table 4, we observe that the proposed SQ-LLaVA achieves significant improvement over the baseline models on the LLaVA (in-the-wild) benchmark by $7.3\%$ and $5.6\%$. This indicates SQ-LLaVA’s capability in more challenging tasks, generalizability to novel domains, and robustness to different prompts.
Table 4: Ablation study of training strategy on visual instruction tasks. All models are in 13B scale with three components of ViT-LoRA (V-LoRA), self-questioning (SQ), and prototype extractor (Proto). We provide experiments on two instruction datasets Liu et al. (2023b); Chen et al. (2023) with different pre-training (PT) and instruction tuning (IT) data scales.
[…]
558K 665K 53.9 68.2 60.2 86.7 69.5 67.7
54.5 70.4 60.6 87.7 71.6 69.0
54.8 69.4 60.7 87.5 72.8 69.0
54.1 68.6 60.1 88.0 71.0 68.4
54.6 69.8 60.2 87.7 74.6 69.4
1200K 700K 53.6 70.1 62.0 87.0 76.4 69.8
56.8 71.0 61.9 87.4 79.6 71.3
54.3 71.1 61.7 86.9 77.0 70.2
56.2 70.2 61.4 86.9 83.6 71.7
58.2 71.7 62.0 87.4 80.7 72.0
A.2 Evaluation on An Additional Benchmark
We evaluate SQ-LLaVA on one recent benchmark MME Fu et al. (2023). MME is designed to assess the performance of vision-language models across a wide range of tasks that require both perception and cognition abilities. In Table 5, we provide a comparison between SQ-LLaVA with other methods. As can be seen, SQ-LLaVA achieves the highest cognition score among all baseline models, indicating a strong performance when handling the commonsense reasoning, numerical calculation, text translation, and code reasoning tasks.
Table 5: Comparison with SOTA methods on recent evaluation benchmarks. The best results are bold and the second-best results are underlined.
[…]
InstructBLIP Dai et al. (2023) 1212.8 291.8
LLaVA-v1.5 Liu et al. (2023b) 1531.3 295.4
ShareGPT4V Chen et al. (2023) 1618.7 303.2
SQ-LLaVA 1568.2 326.8
Appendix B Qualitative Analysis of SQ-LLaVA
B.1 Reliable Detailed Image Description
The detailed image description examples demonstrate that SQ-LLaVA can capture low-level visual information within the image, such as “KL8-Q17” in Fig. 6 and “number 186 206-7” in Fig. 7. Also, SQ-LLaVA demonstrates a better reliability. In Fig. 7, the image shows a complex environment with various objects, which makes it easy for the models to generate wrong descriptions such as “total five people” (LLaVA-v1.5) and “faces blurred” (ShareGPT4V). In contrast, SQ-LLaVA describes uncertainty information with a prefix hedge word such as “likely”, “possibly”, and “perhaps”, achieving higher reliability in a real-world application.
B.2 Visual Self-questioning
In Fig. 10, Fig. 11, and Fig. 12, we show visual self-questioning and answering of SQ-LLaVA. As can be seen, SQ-LLaVA asks meaningful questions based on comprehensive visual understanding. Specifically, the self-asked questions always involve one or more main objects within the image and a relatively divergent problem, where the main objects rely on visual perception, and the meaningful problem requires vision-language alignment. Unlike question-answering with direct instruction, self-questioning encourages the model to discover all possible combinations between visual and semantic information.
Figure 6: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 7: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 8: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 9: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 10: Visual self-questioning and answering of SQ-LLaVA.
Figure 11: Visual self-questioning and answering of SQ-LLaVA.
Figure 12: Visual self-questioning and answering of SQ-LLaVA.
Figure 13: Visual self-questioning and answering of SQ-LLaVA.
Figure 14: Visual self-questioning and answering of SQ-LLaVA.

### v2

@@ 4.5 Ablation Study
Table 3: Ablation study of training strategy on visual instruction tasks. All models are in 7B scale with three components of ViT-LoRA (V-LoRA), self-questioning (SQ), and prototype extractor (Proto). We provide experiments on two instruction datasets [21, 5] with different pre-training (PT) and instruction tuning (IT) data scales.
PT IT V-LoRA SQ Proto VizWiz SQAI VQAT POPE LLaVAW Avg.
@@
5 Conclusions
@@
This work has introduced SQ-LLaVA, a new visual instruction tuning method that enhances general-purpose vision-language understanding and image-oriented question answering through visual self-questioning. Our experiments demonstrate that SQ-LLaVA achieves superior performance with fewer training parameters and instructional data. We have also conducted a comprehensive study on visual discovery/reasoning tasks and found that SQ-LLaVA generalizes well to a wide range of unseen tasks and outperforms several state-of-the-art methods. Qualitative assessments show that SQ-LLaVA strengthens visual representation and domain alignment, effectively reducing object hallucination and improving the semantic interpretation of images. Our findings highlight the potential of visual self-questioning as a powerful training strategy for the visual instruction tuning framework, paving the way for realizing more efficient and effective large vision-language models. Particularly, SQ-LLaVA frames questioning as an intrinsic goal of tuning LLMs, encouraging the exploration of the model’s curiosity (the ability to ask questions proactively) in solving complex problems.